# Airway Labeling Meets Clinical Applications: Reflecting Topology Consistency and Abnormality via Learnable Attentions

### v1

@@ 3.1 Datasets and implementation details
We used a subset of ATM22 [4] for training, selected based on the absence of significant airway deformities, with visibility extending two generations below the subsegmental level. The slice thicknesses were below 0.67 mm, and the spatial resolution ranged from 0.78 mm to 0.82 mm. And 89 cases were selected from the AIIB23 [12] dataset for verifying, exhibiting various airway abnormalities such as atrophy and distortion. Then we extracted the airway centerlines from segmentation results through the skeletonization algorithm [15], identifying bifurcation points to divide the airway into branches. We adopted a three-level nomenclature [16] for the bronchial tree, including 6 types of lobar bronchi, 19 types of segmental bronchi, and 127 types of subsegmental bronchi, with an additional ”outlier” category for abnormal nodes.
The weight of loss function $\gamma_{i}$, $\alpha_{m}$, $\beta_{1}$, and $\beta_{2}$ were set to 1. The label smoothing hyperparameter was set to 0.01. We employed the Adam optimizer for training with a learning rate of $5e^{-4}$ over 600 epochs. For lobar and segmental airway labeling, we stacked two transformers calculated as Eq. (4). At the subsegmental level, we stacked two transformers calculated as Equation 16, followed by two more as Eq. (4). Finally, the predictions from the guided refinement stage served as the output.
@@ 3.2 Evaluation metrics and results
Five evaluation metrics are used to evaluate our method: consistency (CS), accuracy (ACC), precision (PR), recall (RC), and F1 score (F1). The consistency metric was developed to quantify clinical usability. This metric evaluates the anatomical uniformity of predictions for each node and its descendants at the segmental level, defined as the proportion of nodes meeting this criterion.
Table 1: Quantitative comparison with different methods in airway anatomical labeling (%). The best result is in BOLD.
Lobar Segmental Subsegmental
Method ACC PR RC F1 CS ACC PR RC F1 ACC PR RC F1
GCN[17] 91.7 92.0 89.1 89.0 60.4 67.3 59.6 68.7 55.1 45.5 30.5 35.4 29.5
GraphSAGE[18] 93.3 93.6 92.0 91.6 64.4 72.6 64.2 74.0 60.6 55.7 46.0 49.8 44.6
UniSAGE[19] 93.2 93.7 91.9 91.5 69.8 72.9 64.3 74.0 60.6 57.0 46.9 50.6 45.6
TNN[8] 95.3 93.9 94.0 93.2 85.4 78.7 65.2 77.9 64.4 63.3 50.6 57.0 50.9
AirwayFormer[9] 97.1 96.5 97.0 96.6 93.7 90.2 88.9 90.0 88.1 82.3 78.1 79.6 77.5
DeepAssign[20] 96.4 95.9 95.8 95.4 87.1 88.9 87.5 88.4 86.5 81.0 75.3 77.5 74.8
Ours 97.4 96.9 97.2 96.8 96.8 91.4 89.0 90.2 88.5 83.7 79.1 80.9 78.6
@@
Figure 4: Qualitative results at subsegmental label for deformed airways. Ground truth of branches are in black, while false predictions of branches are in red. Our method demonstrates improved accuracy and consistency in predicting deformed airways.
Figure 5: Qualitative results at subsegmental label for abnormal branches. Our method achieves superior performace.
Table 1 reports the quantitative results of our proposed method with other methods. For GNN methods such as GCN[17] and GraphSAGE[18], the one-hot neighborhood aggregation limited their ability to capture long-range topological relationships, resulting in poor classification accuracy, especially at subsegmental level. For HGNN methods, UniSAGE[19] used hyperedges to encode subtree representations. TNN[8] leveraged a novel sub-network architecture for enhanced subtree communication, resulting in improved performance compared to other HGNN methods. For transformer method, the DeepAssign method[20], despite its flexible and differentiable label assignment, faced challenges in learning the complex structure inherent in a large number of subsegmental categories. AirwayFormer[9] leveraged attention mechanisms for global filed and encoded neighborhood information to capture tree structure priors. Compared with the above methods, our method demonstrated a 1.2% increase in segmental level accuracy, a 1.4% increase in subsegmental accuracy, and a remarkable 3.1% improvement in consistency, greatly enhancing clinical value.
Fig. 4 and Fig. 5 qualitatively demonstrate the superiority of our method. For disease-induced airway deformities, the subtrees constructed by the TNN method failed to reflect anatomical meaning, resulting in the misclassification of atrophied regions as the same label. AirwayFormer, which focuses on node-pair relationships without subtree encoding, suffers from inconsistent predictions. For abnormal branches, if they are classified equally with other classes, the absence of a standardized definition for these branches leads to inconsistent predictions. Our method consistently yields accurate predictions in both scenarios. For more details, please refer to the supplementary material.
@@ 3.3 Ablation study
Table 2: Results (%) of ablation study of key components. F2C denotes network with fine-to-coarse guidance.
Lobar Segmental Subsegmental
Method ACC PR RC F1 CS ACC PR RC F1 ACC PR RC F1
Baseline 96.8 96.0 96.4 95.8 87.9 89.2 87.7 88.7 85.7 81.2 76.0 78.1 75.4
F2C 97.2 96.7 97.1 96.6 94.3 90.6 89.3 90.2 88.5 82.7 78.4 80.0 77.9
F2C + SSC 97.4 97.1 97.2 96.8 95.7 90.8 88.2 89.5 87.6 83.2 78.5 80.2 78.0
F2C + ABS 97.5 97.1 97.2 96.9 94.1 91.1 89.6 90.6 90.0 83.4 78.8 80.7 78.6
F2C + SSC + ABS 97.4 96.9 97.2 96.8 96.8 91.4 89.0 90.2 88.5 83.7 79.1 80.9 78.6
Table. 2 reveals the effectiveness of each component in the proposed method. A baseline model consists of three cascaded transformer blocks. Fine-to-coarse guidance encodes the dependencies among the nomenclatures, and their interaction improves the classification results. The incorporation of the SSC module generates a soft subtree map that integrates anatomically relevant topological information, further enhancing performance, particularly a 1.4% increase in consistency. ABS module is effective to suppress the misclassification of the abnormal branches based on the anomaly scores, achieving 83.4% in subsegmental accuracy. The combined modules achieved the best performance, showing improvements of 2.2% in segmental accuracy, 2.5% in subsegmental accuracy, and 8.9% in consistency, thus validating the contribution of each component.
@@ 4 Conclusion
In this paper, an automated method for airway anatomical labeling is proposed. A soft subtree map, incorporating anatomically topological information, was introduced to softly encode subtree representations. Further, it effectively handled anomalous branches through the interaction between node features and prototypes. The superior performance on the AIIB23 dataset, particularly in consistency metrics, demonstrated the clinical potential of our approach.

### v2

@@ 3.1 Datasets and implementation details
We constructed ATL (Airway Tree Labeling) dataset by developing detailed subsegmental anatomical annotations encompassing up to 127 categories, based on chest CT-based binary segmentation mask from ATM22 [5] and AIIB23 [4]. Cases were selected based on visibility extending at least two generations beyond the subsegmental level to facilitate accurate airway labeling. The ATL-clean subset consisting of 196 cases derived from the ATM22 dataset, characterized by the absence of major airway deformities, was used for training and validation. For testing, the ATL-fibrosis subset includes 89 cases derived from AIIB23, a fibrosis airway dataset. This subset encompasses a diverse range of airways, including both normal anatomies and those with extensive deformities, enabling a comprehensive evaluation across both challenging and normal scenarios.
Following the pipeline proposed in TNN[9], airway centerlines were extracted from the binary segmentation results using the skeletonization algorithm [15], and bifurcation points were identified to divide the airway into branches. Branch-wise features were then automatically extracted based on the feature design outlined in TNN[9]. A three level nomenclature [16] was adopted for the bronchial tree, comprising 6 types of lobar bronchi, 19 types of segmental bronchi, and 127 types of subsegmental bronchi, with an additional ”outlier” category for abnormal nodes. Each segment could include three subsegments named ‘a’, ‘b’, and ‘c’ and the branch between bifurcations is named ‘a+b’, ‘b+c’ or ‘a+c’.
The weight of loss function $\gamma_{i}$, $\alpha_{m}$, $\beta_{1}$, and $\beta_{2}$ were set to 1. The label smoothing hyperparameter was set to 0.01. We employed the Adam optimizer for training with a learning rate of $5e^{-4}$ over 600 epochs. For lobar and segmental airway labeling, we stacked two transformers calculated as Eq. (7). At the subsegmental level, we stacked two transformers calculated as Eq. (14), followed by two more as Eq. (7). Finally, the predictions from the guided refinement stage served as the output. Branches with an abnormal score greater than a threshold of 0.1 were predicted as abnormal branches. A sensitivity analysis of this threshold (see Appendix A) confirms its robustness. All experimental results are averaged over five runs with different random seeds to minimize the influence of randomness.
@@ 3.2 Evaluation metrics and results
Four branch-wise evaluation metrics are used: accuracy (ACC), precision (PR), recall (RC), and F1 score (F1). Additionally, two subtree-wise metrics were introduced: Subtree Consistency (SC), which evaluates the proportion of subtrees meeting anatomical labeling criteria, and Topological Distance (TD), which measures the shortest path length between predicted branches and their corresponding ground truth subtrees. The criteria for SC are:
  1. 1.
    Segmental Consistency: All nodes in a subtree share the same predicted segmental label.
  2. 2.
    Subsegmental Consistency: All nodes in a subtree share the same predicted segmental label and conform to the defined subsegmental topology, with nodes labeled as ‘a’, ‘b’, ‘c’, or bifurcation branches (‘a+b’, ‘b+c’, or ‘a+c’).
The evaluation metrics are formally defined as:
$SC=\frac{N_{cs}}{N_{s}}\qquad TD=\frac{1}{N}\sum_{i=1}^{N}\min_{j\in\{{j|y_{j}=\hat{y}_{i}\}}}d_{(i,j)},$ (16)
where $N_{s}$ is the total number of subtrees and $N_{cs}$ denotes the consistent classified number. $N$ is the total number of nodes in a graph, and $d_{i,j}$ denotes the shortest path length between node $v_{i}$ and node $v_{j}$. $y_{j}$ and $\hat{y}_{i}$ denote the ground truth and prediction of node $v_{i}$ respectively.
Table 1: Quantitative comparison with different methods in airway anatomical labeling. PR, RC, and F1 are reported as percentages (%). TD represents a unitless topological distance.
Branch-wise Subtree-wise
Lob Seg Subsegmental Abnormal Segmental Subsegmental
@@
GCN[17] 91.7 67.3 45.5 30.5 35.4 29.5 0.0 0.0 0.0 60.4 2.17 43.9 4.04
GraphSAGE[18] 93.3 72.6 55.7 46.0 49.8 44.6 8.2 2.7 3.6 64.4 1.86 45.3 3.35
UniSAGE[19] 93.2 72.9 57.0 46.9 50.6 45.6 4.7 1.8 2.5 69.8 1.66 62.4 2.97
TNN[9] 95.3 78.7 63.3 50.6 57.0 50.9 1.9 0.9 1.1 85.4 1.31 73.2 2.50
AirwayFormer[10] 97.1 90.2 82.3 78.1 79.6 77.5 45.6 29.0 33.1 93.7 0.64 85.7 1.26
DeepAssign[20] 96.4 88.9 81.0 75.3 77.5 74.8 36.5 17.7 21.1 87.1 0.76 78.7 1.41
Ours 97.4 91.4 83.7 79.1 80.9 78.6 55.6 62.9 54.8 96.8 0.54 86.3 1.17
[…]
11footnotetext: Bold values indicate the best result, and results for ”Ours” are statistically significant ($p\leq 0.05$).
Table 1 presents the quantitative comparison with other methods. GNN methods such as GCN [17] and GraphSAGE [18] suffer from poor classification accuracy, especially at the subsegmental level, due to the limitations of one-hot neighborhood aggregation in capturing long-range topological relationships. HGNN methods, such as UniSAGE [19], utilize hyperedges for subtree representations, while TNN [9] improves subtree communication via a novel sub-network architecture for enhanced subtree communication. Transformer-based methods like DeepAssign [20] demonstrate flexible label assignment but struggle with the complexity of subsegmental categories. AirwayFormer [10] employs attention mechanisms to encode tree structure priors, achieving better performance compared to other baselines. Notably, these methods do not explicitly address abnormal branches and instead treat them as a separate category, limiting their ability to handle the imbalance and undefined nature of these branches effectively.
Compared to AirwayFormer, our method achieves notable improvements. In the branch-wise evaluation, abnormal branch prediction achieves Precision, Recall, and F1 score increases of 10.0%, 33.9%, and 21.9%, respectively. In the subtree-wise evaluation, SC improves by 3.1% at the segmental level and 0.6% at the subsegmental level, while TD decreases by 0.10 and 0.09, respectively. Clinically, these improvements enhance reliability in identifying and navigating intricate airway branches, reducing the risk of errors in bronchoscopic navigation and subsegmentectomy, particularly in cases with abnormal or highly variable branches.
Figure 4: Qualitative results at subsegmental label for deformed airways. Ground truth of branches are in black, while false predictions of branches are in red. TNN fails to generate anatomically meaningful subtrees, while AirwayFormer predicts incorrect labels. Incorporating the SSC module, our method demonstrates improved accuracy and consistency
Figure 5: Qualitative results at subsegmental label for abnormal branches. Competing methods misclassify abnormal branch by treating it as a separate category. Incorporating the ABS module, our method achieves superior performace
Fig. 4 and 5 qualitatively demonstrate our method’s superiority in different challenging scenarios. For disease-induced airway deformities(as shown in Fig. 4), TNN fails to produce anatomically meaningful subtrees, misclassifying atrophied regions. AirwayFormer, focusing on node-pair relationships without subtree encoding, suffers from inconsistent predictions. In contrast, our method leverages the SSC module, which softly integrates topological information to ensure accurate and consistent subtrees. For abnormal branch detection (as shown in Fig. 5), the absence of standardized definitions leads to prediction inconsistencies when treated equally with other categories. In contrast, our method integrates the ABS module, which enables interactions between node features and prototypes to accurately detect abnormal branches while ensuring correct classification of normal branches.
@@ 3.3 Ablation study
Table 2: Results of ablation study of key components. PR, RC, and F1 are reported as percentages (%). TD represents a unitless topological distance. F2C denotes network with fine-to-coarse guidance.
Branch-wise Subtree-wise
Lob Seg Subsegmental Abnormal Segmental Subsegmental
[…]
Baseline 96.8 89.2 81.2 76.0 78.1 75.4 26.0 7.3 9.9 87.9 0.76 78.7 1.41
F2C 97.2 90.6 82.7 78.4 80.0 77.9 44.2 28.7 31.1 94.3 0.64 85.8 1.24
F2C + SSC 97.4 90.8 83.2 78.5 80.2 78.0 43.9 26.9 30.0 95.7 0.61 86.7 1.21
F2C + ABS 97.5 91.1 83.4 78.8 80.7 78.6 59.0 55.1 54.7 94.1 0.58 85.9 1.22
F2C + SSC + ABS 97.4 91.4 83.7 79.1 80.9 78.6 55.6 62.9 54.8 96.8 0.54 86.3 1.17
11footnotetext: Bold values indicate the best result.
Table 2 demonstrates the effectiveness of each component in the proposed method. The baseline model comprises three cascaded Transformer blocks. The fine-to-coarse (F2C) guidance encodes hierarchical dependencies across bronchial nomenclatures, and their interaction improves the classification of both normal and abnormal branches. The SSC module softly encodes subtree representation, resulting in a 1.4% increase in segmental SC and a 0.9% increase in subsegmental SC. The ABS module effectively enhances abnormal branch detection through the interaction between node features and prototypes, achieving 62.9% abnormal branch recall. Additionally, it also improves the classification of normal branches by suppressing the interference of abnormal branches. Overall, the proposed model demonstrates enhanced performance on both branch-wise and subtree-wise evaluation metrics due to the complementary contributions of its individual components.
@@ 4 Conclusion
In this paper, an automated method for airway anatomical labeling is proposed. A soft subtree map, incorporating anatomically topological information, was introduced to softly encode subtree representations. Further, it effectively handled anomalous branches through interactions between node features and prototypes. The superior performance on a challenging dataset highlighting its potential for enhancing topology consistency and detecting abnormalities in clinical scenarios.
Declarations
  • Funding This work was supported in part by National Key R&D Program of China (Grant Number: 2022ZD0212400), Natural Science Foundation of China (Grant Number: 62373243) and the Science and Technology Commission of Shanghai Municipality, China (Grant Number: 20DZ2220400), Shanghai Municipal Science and Technology Major Project (No. 2021SHZDZX0102).
  • Competing interests The authors have no competing interests to declare that are relevant to the content of this article.
  • Ethics approval This article does not contain any studies with human participants or animals performed by any of the authors.
  • Consent for publication The study conducted experiments based on two publicly available datasets. The access of our airway anatomical annotation can be applied for research purpose upon reasonable request.
Appendix A: Sensitivity Analysis of Abnormal Score Threshold
The appendix presents a sensitivity analysis of the abnormal score threshold in the ABS module. The analysis evaluates precision (PR), recall (RC), and F1 score across abnormal score thresholds ranging from 0.1 to 0.9, as illustrated in Fig. 6. Experiments were conducted on the ATL dataset using our method.
The result demonstrates the flexibility and reliability of ABS module under varying conditions. A threshold of 0.1 achieves the highest F1 score, balancing precision and recall. Notably, our method consistently outperforms AirwayFormer(as shown in Table 1) across all metrics, particularly in Recall and F1 Score, demonstrating its superiority in abnormal branch detection.
Figure 6: The sensitivity analysis of abnormal branch detection performance with varying abnormal score thresholds. Metrics include Precision (PR), Recall (RC), and F1 Score. A branch is classified as abnormal if its abnormal score exceeds the threshold